A company invested in an automated sorting machine for $650,000. The installation costs incurred were $150,000. The CNC mill has a MACRS-GDS 5-year property class. The BT&LCF is $350,000 per year, and the company’s tax rate is 40%.

  1. The company used the all allowable depreciation deductions during the first year. The company has decided to sell the asset during year 4, what is the depreciation allowance for year 1.
Answer #1

Cost basis = 650000 + 150000 = 800000

MACRS GDS rate for depreciation for 5 yrs property are 0.2, 0.32, 0.192, 0.1152, 0.1152 & 0.0576

Depreciation allowance for first yr = 0.2 * 800000 = 160000